# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 11. Regular tree systems of spaces with doubleton peripherals
be a connecting system for regular spaces with doubleton peripherals. For each $j\in J$, denote by $d_{j}:\Omega_{j}\to Z$ the degree function of the extra-decoration ${\mathcal{D}}_{j}^{+}$. We say that $\Theta$ is modelled on $\mathcal{P}$ if each constituent space $L_{t}$ of $\Theta$ can be equipped with an identification (as a space with peripherals) with one of the spaces $K_{j}$ of $\mathcal{P}$ (and hence also equipped with the corresponding extra-decoration ${\mathcal{D}}_{j}^{+}$) so that the following conditions are satisfied:
@@
    for each white vertex $u$ of $T$ there is a block $B$ of ${\mathcal{B}}$ such that the set of types of the peripherals of form $\varphi_{[u,t]}(\Sigma_{u})\subset L_{t}$ (under the identification of $L_{t}$ with one of $K_{j}$’s, as above), where $t$ runs through the set of all vertices of $T$ adjacent to $u$, coincides with $B$;
@@
    for any $u$ and $B$ as in (3) (i.e. with $B$ non-orientable), if $\lambda\in B$ is an orientable type then, again denoting by $K_{j}$ this unique space of $\mathcal{P}$ which contains peripherals of type $\lambda$, the number of (black) vertices $t$ adjacent to $u$ for which the spaces $L_{t}$ are identified with $K_{j}$ and the peripherals $\varphi_{[u,t]}(\Sigma_{u})$ are of type $\lambda$ is equal to twice the degree, i.e. to $2\cdot d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ (for any $t$ as above); moreover, if we choose an auxiliary orientation of the doubleton $\Sigma_{u}$ (i.e. an order for its two points) then for exactly half of the vertices $t$ as above (i.e. for $d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ of those vertices) the map $\varphi_{[u,t]}$ preserves the orientations, and for another half of such vertices it reverses the orientations.
@@
Next example will play a central role in the final step of the proof of Theorem 10.1, as given in Section 13.
@@
We will show that any tree system $\Theta^{r}(G)$, as described in Subsection 3.2, has a natural structure of a regular tree system of spaces with doubleton peripherals. We will do this by explicitely describing a connecting system ${\mathcal{P}}={\mathcal{P}}^{r}_{G}$ on which the corresponding tree system $\Theta^{r}(G)$ is modelled. We refer to the terminology and notation as in Subsection 3.2.
@@
For any $(K_{j},\Omega_{j})$ as above, take as a decoration ${\mathcal{D}}_{j}$ all the data as in Example 11.2, under viewing $G(v_{j})$ as an abstract factor, with the orientations of orientable peripheral subgroups chosen in a way described in the previous paragraph. To describe the appropriate degree function $d_{j}$, consider any peripheral $\partial G(u)$ in $K_{j}=\partial G(v_{j})$, where $u$ is some (white) vertex of $T^{r}_{G}$ adjacent to $v_{j}$. Recall that the peripheral subgroup corresponding to this peripheral of $K_{j}$ is $G([u,v_{j}])=G(u)\cap G(v_{j})$, which we denote shortly as $H_{u}$. If $G(u)$ (which is a 2-ended group) is orientable, we take as $d_{j}(\partial G(u))$ the index $[G(u):H_{u}]$. If both $G(u)$ and $H_{u}$ are non-orientable, we similarly put $d_{j}(\partial G(u)):=[G(u):H_{u}]$. However, if $G(u)$ is non-orientable, and $H_{u}$ is orientable, we notice that the index $[G(u):H_{u}]$ is an even number, and we take as $d_{j}(\partial G(u))$ half of this number. Note that, due to the final comment in Example 11.2, with the so described extra-decoration ${\mathcal{D}}_{j}^{+}=({\mathcal{D}}_{j},d_{j})$, each of the spaces $(K_{j},\Omega_{j})$ becomes an extra-decorated regular space with doubleton peripherals.
We now pass to the description of an appropriate data =${\mathcal{B}}=\{B_{a}:a\in A\}$. Take as $A$ the set of $G$-orbits of the white vertices of $T^{r}_{G}$. For any orbit $a\in A$, let $u$ be any vertex in this orbit. Note that for any (black) vertex $t$ adjacent to $u$, the corresponding space with peripherals $\partial G(t)$ is identified, uniquely up to a decoration preserving homeomorphism, with the appropriate space $K_{j}=\partial G(v_{j})$, where $v_{j}$ is in the same orbit as $t$, via any of the maps from the $G$-action of $G$ on $\partial G$ which sends $\partial G(v_{j})$ to $\partial G(t)$. In particular, under the latter identifications, all peripherals in all such spaces $\partial G(t)$ have well defined types. Define $B_{a}$ to be the set of all types appearing among the peripherals of the form $\varphi_{[u,t]}(\Sigma_{u})$ (in the corresponding spaces $\partial G(t)$), where $t$ runs through all (black) vertices of $T^{r}_{G}$ adjacent to $u$. Furthermore, we declare that the block $B_{a}$ is orientable if and only if the corresponding 2-ended group $G(u)$ is orientable.
@@
    ${\mathcal{T}}_{j}=\{[e]\in\mathcal{T}_{G}:\hbox{ $e$ is adjacent to $v_{j}$}\},$
@@
      if both groups $G(e),G(u)$ ore orientable, or both are non-orientable, the degree $d_{j}([e])$ coincides with the number of edges adjacent to $u$ in the $G$-orbit of $e$;
@@
Note that, under notation of the above Example 11.10, if $v_{j}$ is a flexible vertex, the decorated space with doubleton peripherals $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ is a densely decorated punctured circle, as described in Example 11.3 (for justification, see Example 8.8 and the last paragraph in Example 11.3). This densely decorated punctured circle is easily seen to be orientable if and only if the flexible factor corresponding to the group $G(v_{j})$ is orientable. The numbers of types of orientable and non-orientable peripherals in $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ coincide with the numbers of conjugacy classes in $G(v_{j})$ among the orientable (and non-orientable, respectively) subgroups of the form $G(u)\cap G(v_{j}):u\in N^{T^{r}_{G}}_{v_{j}}$. As we have already noticed before, the informations from the last two sentences determine the decorated space $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ uniquely up to an isomorphism (see Lemma 11.4(2)(3)).
@@ 12. Trees of internally punctured segments
[…]
[N: This note remained from some time ago: “Need to include some discussion of trees of punctured circles.” We should keep this in mind as we read this section: is there adequate discussion of trees of punctured circles/segments.]
@@ 12.1. Orientable trees of segments
    identifying naturally $I^{\circ}_{\rm int}$ with the punctured circle $(S^{1})^{\circ}$, and fixing one of its orientations (in the sense of the last sentence of Definition 8.10), we equip all its peripherals (both ordinary and special ones) with the orientations compatible with the just fixed orientation of the whole space;
@@
Define the tree system $\Psi^{o}_{m}(I^{\circ}_{\rm int})$ by requiring that:
@@
    its constituent spaces are the copies of the internally punctured segments $I_{\tau}^{\circ}:\tau\in{\mathcal{T}}$, with ordinary peripherals as the peripherals of these spaces (it means that the special peripherls are not among the peripherals of those constituent spaces);
@@
    gluing maps respect the orientations of the glued peripherals (which means that those maps reverse the orientations of the glued peripherals);
@@
Consider the limit $X$ of the tree system $\Psi^{o}_{m}(I^{\circ}_{\rm int})$ of punctured segments, and view it as a space with peripherals, where peripherals are the images of all special peripherals in all constituent punctured segments. Equip also all those peripherals in a natural way with the types from the set $\mathcal{T}$, and with the naturally induced orientations. Then $X$ is homeomorphic, as a space with peripherals, to an orientable densely decorated punctured circle, as described in Example 11.3, and it has precisely $m$ types of the peripherals.
@@
At first, we will ignore the types and the orientations of the peripherals, and show that, as a pure space with peripherals, the limit of the tree system $\Psi^{o}_{m}(I^{\circ}_{\rm int})$ is homeomorphic to the punctured circle $(S^{1})^{\circ}$. To do this, consider a related tree system $\overline{\Psi}^{o}_{m}(I^{\circ}_{\rm int})$ with the same underlying tree $T$, in which the constituent spaces $\overline{K}_{t}$ are obtained from the corresponding constituent spaces $K_{t}$ of $\Psi^{o}_{m}(I^{\circ}_{\rm int})$ (which are punctured segments) as follows. Glue to each $K_{t}$ a segment $I_{t}$ via a gluing map which sends the special peripheral of $K_{t}$ bijectively to the boundary $\partial I_{t}$. Identify the (ordinary) peripherals in $K_{t}$ as subspaces of $\overline{K}_{t}$, and view them as the peripherals of $\overline{K}_{t}$. For the gluing maps (or more formally, connecting maps) in the new system $\overline{\Psi}^{o}_{m}(I^{\circ}_{\rm int})$ take then the same maps which appeared as connecting maps in the system $\Psi^{o}_{m}(I^{\circ}_{\rm int})$.
@@
Note that the larger constituent spaces $\overline{K}_{t}$ are not homeomorphic (as spaces with peripherals) to the punctured circle, or to any other punctured graph. However, one can still associate to the tree system $\overline{\Psi}:=\overline{\Psi}^{o}_{m}(I^{\circ}_{\rm int})$, in a way similar as in Subsection 2.6, the inverse system ${\mathcal{S}}_{\overline{\Psi}}$, and by the arguments similar as those presented in Subsection 2.6, this inverse system has the property that $\lim_{\leftarrow}{\mathcal{S}}_{\overline{\Psi}}\cong\lim\overline{\Psi}$. Now, since all the spaces in the inverse system ${\mathcal{S}}_{\overline{\Psi}}$ are clearly homeomorphic to the circle $S^{1}$, and since all bonding maps in this system consist of shrinking finitely many pairwise disjoint closed subsegments of $S^{1}$ to points, these maps are obviously near homeomorphisms, i.e. they can be approximated by homeomorphisms. Since, by a result of M. Brown ([3, Theorem 4]), the inverse limit maps in the inverse system consisting of near homeomorphisms are near homeomorphisms, it follows that the inverse limit $\lim_{\leftarrow}{\mathcal{S}}_{\overline{\Psi}}$ is homeomorphic to $S^{1}$, and thus we also have $\lim\overline{\Psi}\cong S^{1}$.
@@
The segments $I_{t}$ from the constituent spaces $\overline{K}_{t}$ obviously become a subsegments in the limit circle $\lim\overline{\Psi}$, and moreover they are easily seen to be pairwise disjoint. A slightly less clear fact, the straightforward proof of which we omit, is that the union of the segments $I_{t}$ (viewed as subsegments in the circle $\lim\overline{\Psi}$) is dense in $\lim\overline{\Psi}$. On the other hand, the limit $\lim\Psi$ of the original tree system $\Psi=\Psi^{o}_{m}(I^{\circ}_{\rm int})$ can be identified with a subspace in the limit $\lim\overline{\Psi}$, namely the subspace obtained by deleting from $\lim\overline{\Psi}$ the interiors of all subsegments $I_{t}$. Moreover, the peripherals in $\lim\Psi$ correspond to the pairs of endpoints of the deleted subsegments $I_{t}$. Consequently, as a space with peripherals, $\lim\Psi$ is homeomorphic to the punctured circle, as required.
@@
It remains to check the conditions related to the types and orientations of the peripherals. Condition which requires that the orientations of all peripherals in the limit space $\lim\Psi$ are consistent (as in condition (b) in Example 11.3) follows directly from condition (c) above in this subsection. Condition which requires that the union of the peripherals of any fixed type is dense follows easily from condition (d) above. This completes the proof. ∎
@@ 12.2. Non-orientable trees of segments
    for each type $\tau\in{\mathcal{T}}_{1}$, we equip the special peripheral of $I^{\circ}_{\tau}$ with one of the orientations, and for each type $\tau\in{\mathcal{T}}_{2}$ we associate to the special peripheral of $I^{\circ}_{\tau}$ its both orientations.
Define the tree system $\Psi^{n-o}_{m,n}(I^{\circ}_{\rm int})$ by requiring that:
@@
    for each type $\tau\in{\mathcal{T}}$, and each type $\tau^{\prime}\in{\mathcal{T}}_{1}$, any constituent space $K_{t}$ being a copy of $I^{\circ}_{\tau}$, after fixing its auxiliary orientation, satisfies the following condition:
@@
Consider the limit $X$ of the tree system $\Psi^{n-o}_{m,n}(I^{\circ}_{\rm int})$ of punctured segments, and view it as a space with peripherals, where peripherals are the images of all special peripherals in all constituent punctured segments. Equip also all those peripherals in a natural way with the types from the set $\mathcal{T}$, and with the naturally induced orientations (one orientation if $\tau\in{\mathcal{T}}_{1}$, and both orientations if $\tau\in{\mathcal{T}}_{2}$). Then $X$ is homeomorphic, as a space with peripherals, to a non-orientable densely decorated punctured circle, as described in Example 11.3, and it has precisely $m$ types of orientable peripherals, and $n$ types of non-orientable peripherals.
@@
The major part of this section is devoted to the proof of Theorem 10.1. After doing this, we provide also arguments for getting (as rather immediate corollaries to Theorem 10.1) the proofs of the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B.
The outline of the proof of Theorem 10.1 is as follows. Given a group $G$ as in the assumption of Theorem 10.1, we have described two tree systems canonically associated to $G$. The first one is the tree system $\Theta^{r}(G)$ associated to the reduced Bowditch JSJ splitting of $G$, as described in Subsection 3.2, and it has the property that $\lim\Theta^{r}(G)\cong\partial G$ (see Theorem 3.8). The second tree system associated to $G$ is the (regular) tree system of punctured graphs $\Theta[{\mathcal{R}}_{G}]$ associated to the graphical connecting system ${\mathcal{R}}_{G}$, as described in Section 4 (together with the description of ${\mathcal{R}}_{G}$ given in Section 10), and as appearing in the statement of Theorem 10.1. In view of the equality $\partial G\cong\lim\Theta^{r}(G)$, our goal is to show that $\lim\Theta^{r}(G)\cong\lim\Theta[{\mathcal{R}}_{G}]$.
In order to show the latter equality, we shall describe three more tree systems, denoted $\Theta^{V}[{\mathcal{R}}_{G}]$, $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ and $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, with the following properties. The tree system $\Theta^{V}[{\mathcal{R}}_{G}]$, called the tree system of V-trees associated to $G$, will be described as obtained from $\Theta[{\mathcal{R}}_{G}]$ via certain consolidation (an operation described in Subsection 2.7, which does not affect the limit space of the system). The constituent spaces of $\Theta^{V}[{\mathcal{R}}_{G}]$ will be the V-trees of graphs given as limits of certain V-tree systems of graphs naturally appearing inside the system $\Theta[{\mathcal{R}}_{G}]$. The tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, called the tree system of cores and arms associated to $G$, will be described as obtained from $\Theta^{V}[{\mathcal{R}}_{G}]$ by an operation of splitting (i.e. operation opposite to some operation of consolidation) consisting of appropriate splittings of all V-tree constituent spaces of $\Theta^{V}[{\mathcal{R}}_{G}]$ into cores and arms (as described in Section 7). Finally, the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, called the tree system of cores and consolidated arms associated to $G$, will be described as obtained from $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ via one more operation of consolidation, which will join together certain configurations of arms in the latter system into bigger constituent spaces which have the form of punctured circles (here we will refer to the results of Section 12).

### v2

@@ 11. Regular tree systems of spaces with doubleton peripherals
be a connecting system for regular spaces with doubleton peripherals. For each $j\in J$, denote by $d_{j}:\Omega_{j}\to\mathbb{Z}$ the degree function of the extra-decoration ${\mathcal{D}}_{j}^{+}$. We say that $\Theta$ is modelled on $\mathcal{P}$ if each constituent space $L_{t}$ of $\Theta$ can be equipped with an identification (as a space with peripherals) with one of the spaces $K_{j}$ of $\mathcal{P}$ (and hence also equipped with the corresponding extra-decoration ${\mathcal{D}}_{j}^{+}$) so that the following conditions are satisfied:
@@
    for each white vertex $u$ of $T$ there is a block $B$ of ${\mathcal{B}}$ such that the set of types of the peripherals of the form $\varphi_{[u,t]}(\Sigma_{u})\subset L_{t}$ (under the identification of $L_{t}$ with one of the $K_{j}$, as above), where $t$ runs through the set of all vertices of $T$ adjacent to $u$, coincides with $B$;
@@
    for any $u$ and $B$ as in (3) (i.e. with $B$ non-orientable), if $\lambda\in B$ is an orientable type then, again denoting by $K_{j}$ this unique space of $\mathcal{P}$ which contains peripherals of type $\lambda$, the number of (black) vertices $t$ adjacent to $u$ for which the spaces $L_{t}$ are identified with $K_{j}$ and the peripherals $\varphi_{[u,t]}(\Sigma_{u})$ are of type $\lambda$ is equal to twice the degree, i.e., to $2\cdot d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ (for any $t$ as above); moreover, if we choose an auxiliary orientation of the doubleton $\Sigma_{u}$ (i.e. an order for its two points) then for exactly half of the vertices $t$ as above (i.e. for $d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ of those vertices) the map $\varphi_{[u,t]}$ preserves the orientations, and for another half of such vertices it reverses the orientations.
@@
The next example will play a central role in the final step of the proof of Theorem 10.1, as given in Section 13.
@@
We will show that any tree system $\Theta^{r}(G)$, as described in Section 3.2, has a natural structure of a regular tree system of spaces with doubleton peripherals. We will do this by explicitly describing a connecting system ${\mathcal{P}}={\mathcal{P}}^{r}_{G}$ on which the corresponding tree system $\Theta^{r}(G)$ is modelled. We refer to the terminology and notation as in Section 3.2.
@@
For any $(K_{j},\Omega_{j})$ as above, take as a decoration ${\mathcal{D}}_{j}$ all the data as in Example 11.2, under viewing $G(v_{j})$ as an abstract factor, with the orientations of orientable peripheral subgroups chosen in a way described in the previous paragraph. To describe the appropriate degree function $d_{j}$, consider any peripheral $\partial G(u)$ in $K_{j}=\partial G(v_{j})$, where $u$ is some (white) vertex of $T^{r}_{G}$ adjacent to $v_{j}$. Recall that the peripheral subgroup corresponding to this peripheral of $K_{j}$ is $G([u,v_{j}])=G(u)\cap G(v_{j})$, which we denote briefly by $H_{u}$. If $G(u)$ (which is a 2-ended group) is orientable, we take as $d_{j}(\partial G(u))$ the index $[G(u):H_{u}]$. If both $G(u)$ and $H_{u}$ are non-orientable, we similarly put $d_{j}(\partial G(u)):=[G(u):H_{u}]$. However, if $G(u)$ is non-orientable, and $H_{u}$ is orientable, we notice that the index $[G(u):H_{u}]$ is an even number, and we take as $d_{j}(\partial G(u))$ half of this number. Note that, due to the final comment in Example 11.2, with the so described extra-decoration ${\mathcal{D}}_{j}^{+}=({\mathcal{D}}_{j},d_{j})$, each of the spaces $(K_{j},\Omega_{j})$ becomes an extra-decorated regular space with doubleton peripherals.
We now pass to the description of appropriate data =${\mathcal{B}}=\{B_{a}:a\in A\}$. Take as $A$ the set of $G$-orbits of the white vertices of $T^{r}_{G}$. For any orbit $a\in A$, let $u$ be any vertex in this orbit. Note that for any (black) vertex $t$ adjacent to $u$, the corresponding space with peripherals $\partial G(t)$ is identified, uniquely up to a decoration preserving homeomorphism, with the appropriate space $K_{j}=\partial G(v_{j})$, where $v_{j}$ is in the same orbit as $t$, via any of the maps from the $G$-action of $G$ on $\partial G$ which sends $\partial G(v_{j})$ to $\partial G(t)$. In particular, under the latter identifications, all peripherals in all such spaces $\partial G(t)$ have well defined types. Define $B_{a}$ to be the set of all types appearing among the peripherals of the form $\varphi_{[u,t]}(\Sigma_{u})$ (in the corresponding spaces $\partial G(t)$), where $t$ runs through all (black) vertices of $T^{r}_{G}$ adjacent to $u$. Furthermore, we declare that the block $B_{a}$ is orientable if and only if the corresponding 2-ended group $G(u)$ is orientable.
@@
    ${\mathcal{T}}_{j}=\{[e]:\text{ $e$ is an edge of $T^{r}_{G}$ adjacent to $v_{j}$}\},$
@@
      if both groups $G(e),G(u)$ are orientable, or both are non-orientable, the degree $d_{j}([e])$ coincides with the number of edges adjacent to $u$ in the $G$-orbit of $e$;
@@
Note that, under the notation of the above Example 11.10, if $v_{j}$ is a flexible vertex, the decorated space with doubleton peripherals $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ is a densely decorated punctured circle, as described in Example 11.3 (for justification, see Example 8.8 and the last paragraph in Example 11.3). This densely decorated punctured circle is easily seen to be orientable if and only if the flexible factor corresponding to the group $G(v_{j})$ is orientable. The numbers of types of orientable and non-orientable peripherals in $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ coincide with the numbers of conjugacy classes in $G(v_{j})$ among the orientable (and non-orientable, respectively) subgroups of the form $G(u)\cap G(v_{j}):u\in N^{T^{r}_{G}}_{v_{j}}$. As we have already noticed before, the information from the last two sentences determines the decorated space $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ uniquely up to an isomorphism (see Lemma 11.4(2)(3)).
@@ 12.1. Orientable trees of segments
    identifying naturally $I^{\circ}_{\rm int}$ with the punctured circle $(S^{1})^{\circ}$, and fixing one of its orientations (in the sense of the last sentence of Definition 8.10), we equip all its peripherals (both ordinary and special) with the orientations compatible with the just fixed orientation of the whole space;
@@
Define the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ by requiring the following (the superscript “o” on $\Psi^{\text{o}}_{m}$ indicates “orientable”):
@@
    its constituent spaces are the copies of the internally punctured segments $I_{\tau}^{\circ}:\tau\in{\mathcal{T}}$, with ordinary peripherals as the peripherals of these spaces (this means that the special peripherals are not among the peripherals of those constituent spaces);
@@
    gluing maps respect the orientations of the glued peripherals, which means that those maps reverse the orientations of the glued peripherals;
@@
The fact that such a tree system is unique up to an isomorphism follows by discussion and arguments similar to those in Example 11.3, Lemma 11.4 and Lemma 11.9, but this fact is not essential for our further considerations.
@@
Consider the limit $X$ of the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ of punctured segments, and view it as a space with peripherals, where peripherals are the images of all special peripherals in all constituent punctured segments. Equip also all those peripherals in a natural way with the types from the set $\mathcal{T}$, and with the naturally induced orientations. Then $X$ is homeomorphic, as a space with peripherals, to an orientable densely decorated punctured circle, as described in Example 11.3, and it has precisely $m$ types of peripherals.
@@
At first, we will ignore the types and the orientations of the peripherals, and show that, as a pure space with peripherals, the limit of the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ is homeomorphic to the punctured circle $(S^{1})^{\circ}$. To do this, consider a related tree system $\overline{\Psi}^{\text{o}}_{m}(I^{\circ}_{\rm int})$ with the same underlying tree $T$, in which the constituent spaces $\overline{K}_{t}$ are obtained from the corresponding constituent spaces $K_{t}$ of $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ (which are punctured segments) as follows. Glue to each $K_{t}$ a segment $I_{t}$ via a gluing map which sends the special peripheral of $K_{t}$ bijectively to the boundary $\partial I_{t}$. Identify the (ordinary) peripherals in $K_{t}$ as subspaces of $\overline{K}_{t}$, and view them as the peripherals of $\overline{K}_{t}$. For the gluing maps (or more formally, connecting maps) in the new system $\overline{\Psi}^{\text{o}}_{m}(I^{\circ}_{\rm int})$ take then the same maps which appeared as connecting maps in the system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$.
@@
As in Section 2.6, we associate to the tree system $\overline{\Psi}:=\overline{\Psi}^{\text{o}}_{m}(I^{\circ}_{\rm int})$ the inverse system ${\mathcal{S}}_{\overline{\Psi}}$ such that, by Proposition 2.8, $\lim_{\leftarrow}{\mathcal{S}}_{\overline{\Psi}}\cong\lim\overline{\Psi}$. Now, since all the spaces in the inverse system ${\mathcal{S}}_{\overline{\Psi}}$ are clearly homeomorphic to the circle $S^{1}$, and since all bonding maps in this system consist of shrinking finitely many pairwise disjoint closed subsegments of $S^{1}$ to points, these maps are obviously near homeomorphisms, i.e., they can be approximated by homeomorphisms. Since, by a result of M. Brown ([Bro60, Theorem 4]), the inverse limit maps in an inverse system consisting of near homeomorphisms are near homeomorphisms, it follows that the inverse limit $\lim_{\leftarrow}{\mathcal{S}}_{\overline{\Psi}}$ is homeomorphic to $S^{1}$, and thus we also have $\lim\overline{\Psi}\cong S^{1}$.
@@
The segments $I_{t}$ from the constituent spaces $\overline{K}_{t}$ obviously become subsegments in the limit circle $\lim\overline{\Psi}$, and moreover they are easily seen to be pairwise disjoint. A slightly less clear fact, the straightforward proof of which we omit, is that the union of the segments $I_{t}$ (viewed as subsegments in the circle $\lim\overline{\Psi}$) is dense in $\lim\overline{\Psi}$. On the other hand, the limit $\lim\Psi$ of the original tree system $\Psi=\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ can be identified with a subspace in the limit $\lim\overline{\Psi}$, namely the subspace obtained by deleting from $\lim\overline{\Psi}$ the interiors of all subsegments $I_{t}$. Moreover, the peripherals in $\lim\Psi$ correspond to the pairs of endpoints of the deleted subsegments $I_{t}$. Consequently, as a space with peripherals, $\lim\Psi$ is homeomorphic to the punctured circle, as required.
@@
It remains to check the conditions related to the types and orientations of the peripherals. The condition which requires that the orientations of all peripherals in the limit space $\lim\Psi$ are consistent (as in condition (b) in Example 11.3) follows directly from condition (c) above in this subsection. The condition which requires that the union of the peripherals of any fixed type is dense follows easily from condition (d) above. This completes the proof. ∎
@@ 12.2. Non-orientable trees of segments
    for each type $\tau\in{\mathcal{T}}_{1}$, we equip the special peripheral of $I^{\circ}_{\tau}$ with one of the orientations, and for each type $\tau\in{\mathcal{T}}_{2}$ we associate to the special peripheral of $I^{\circ}_{\tau}$ both of its orientations.
Define the tree system $\Psi^{\text{n-o}}_{m,n}(I^{\circ}_{\rm int})$ by requiring the following (the superscript “n-o” on $\Psi^{\text{n-o}}_{m,n}$ indicates “non-orientable”):
@@
    for each type $\tau\in{\mathcal{T}}$, and each type $\tau^{\prime}\in{\mathcal{T}}_{1}$, any constituent space $K_{t}$ being a copy of $I^{\circ}_{\tau}$, after fixing an auxiliary orientation, satisfies the following condition:
@@
Similarly as we have mentioned in the previous subsection, the fact that a tree system as above is unique up to an isomorphism follows by discussion and arguments similar to those in Example 11.3, Lemma 11.4 and Lemma 11.9. However, this fact is not essential for our further considerations.
@@
Consider the limit $X$ of the tree system $\Psi^{\text{n-o}}_{m,n}(I^{\circ}_{\rm int})$ of punctured segments, and view it as a space with peripherals, where peripherals are the images of all special peripherals in all constituent punctured segments. Equip also all those peripherals in a natural way with the types from the set $\mathcal{T}$, and with the naturally induced orientations (one orientation if $\tau\in{\mathcal{T}}_{1}$, and both orientations if $\tau\in{\mathcal{T}}_{2}$). Then $X$ is homeomorphic, as a space with peripherals, to a non-orientable densely decorated punctured circle, as described in Example 11.3, and it has precisely $m$ types of orientable peripherals, and $n$ types of non-orientable peripherals.
@@
The major part of this section is devoted to the proof of Theorem 10.1. After doing this, we also provide arguments for getting (as rather immediate corollaries to Theorem 10.1) the proofs of the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B.
The outline of the proof of Theorem 10.1 is as follows. Given a group $G$ as in the assumption of Theorem 10.1, we have described two tree systems canonically associated to $G$. The first one is the tree system $\Theta^{r}(G)$ associated to the reduced Bowditch JSJ splitting of $G$, as described in Section 3.2, and it has the property that $\lim\Theta^{r}(G)\cong\partial G$ (see Theorem 3.8). The second tree system associated to $G$ is the (regular) tree system of punctured graphs $\Theta[{\mathcal{R}}_{G}]$ associated to the graphical connecting system ${\mathcal{R}}_{G}$, as described in Section 4 (together with the description of ${\mathcal{R}}_{G}$ given in Section 10), and as appearing in the statement of Theorem 10.1. In view of the equality $\partial G\cong\lim\Theta^{r}(G)$, our goal is to show that $\lim\Theta^{r}(G)\cong\lim\Theta[{\mathcal{R}}_{G}]$.
In order to show the latter equality, we shall describe three more tree systems, denoted $\Theta^{V}[{\mathcal{R}}_{G}]$, $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ and $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, with the following properties. The tree system $\Theta^{V}[{\mathcal{R}}_{G}]$, called the tree system of V-trees associated to $G$, will be described as obtained from $\Theta[{\mathcal{R}}_{G}]$ via a certain consolidation (an operation described in Section 2.7, which does not affect the limit space of the system). The constituent spaces of $\Theta^{V}[{\mathcal{R}}_{G}]$ will be the V-trees of graphs given as limits of certain V-tree systems of graphs naturally appearing inside the system $\Theta[{\mathcal{R}}_{G}]$. The tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, called the tree system of cores and arms associated to $G$, will be described as obtained from $\Theta^{V}[{\mathcal{R}}_{G}]$ by an operation of splitting (i.e. operation opposite to some operation of consolidation) consisting of appropriate splittings of all V-tree constituent spaces of $\Theta^{V}[{\mathcal{R}}_{G}]$ into cores and arms (as described in Section 7). Finally, the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, called the tree system of cores and consolidated arms associated to $G$, will be described as obtained from $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ via one more operation of consolidation, which will join together certain configurations of arms in the latter system into bigger constituent spaces which have the form of punctured circles (here we will refer to the results of Section 12).